OpenStudy (anonymous):

Write a polynomial function with rational coefficients so that P(x)=0 has the given roots. -9 and -15

OpenStudy (anonymous):

I think I figured it out! (x+9)(x+15) x^2+24x+135

OpenStudy (jdoe0001):

\(\bf -9, -15\quad meaning\\ \quad \\ x=-9\implies x+9=0\implies (x+9)=0\\ \quad \\ x=-15\implies x+15=0\implies (x+15)=0\\ \quad \\ P(x)=0\implies (x+9)(x+15)=0\)

OpenStudy (jdoe0001):

yeap :)

OpenStudy (anonymous):

Thank you! I understood it as soon as I posted this question. :)